answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What is the probability of drawing 2 cards that are a 2 from a deck of 52 when first card is an ace and 2nd non ace?

The probability of drawing two cards that are a 2 when the first card is an Ace is zero, because a two is not an Ace. They are mutually exclusive events.

If this is not clear, consider the probability of rolling a seven on one roll of one six sided die. That probability is also zero - it will not happen.

What is not dependent on the size of a sample?

In general the mean of a truly random sample is not dependent on the size of a sample. By inference, then, so is the variance and the standard deviation.

What is the probability of getting 6 heads in tossing 6 coins in 6 tosses?

The probability of tossing 6 heads in 6 dice is 1 in 26, or 1 in 64, or 0.015625. THe probability of doing that at least once in six trials, then, is 6 in 26, or 6 in 64, or 3 in 32, or 0.09375.

What is the probability of having two aces next to each other ANYWHERE in a deck of cards?

There are 4 aces in a deck, a deck consists of 52 cards

4/52*3/51 ≈ 0.45% that 2 cards with the same value would end up next to each other

What word can you make with t q t h r u t?

According to the Franklin SCR-226 Scrabble Player's Dictionary (2003), the words that can be formed from the letters TQTHRUT are TRUTH, HURT, RUTH, THRU, HUT, UH, RUT, TUT, and UT. I know of no more. Anyone else???

What is the probability of obtaining a sum of at least 7 when rolling a pair of dice?

There are 36 permutations of two dice. Of these, 21 have a sum of at least 7. The probability, then, of rolling at least a sum of 7 is 21 in 36, or about 0.5833.

The permutation list is below. The sums at least 7 are bold. Count them, and understand.

1+1 = 2

1+2 = 3

1+3 = 4

1+4 = 5

1+5 = 6

1+6 = 7

2+1 = 3

2+2 = 4

2+3 = 5

2+4 = 6

2+5 = 7

2+6 = 8

3+1 = 4

3+2 = 5

3+3 = 6

3+4 = 7

3+5 = 8

3+6 = 9

4+1 = 5

4+2 = 6

4+3 = 7

4+4 = 8

4+5 = 9

4+6 = 10

5+1 = 6

5+2 = 7

5+3 = 8

5+4 = 9

5+5 = 10

5+6 = 11

6+1 = 7

6+2 = 8

6+3 = 9

6+4 = 10

6+5 = 11

6+6 = 12

How many standard deviations above or below the mean should a score be to establish significance?

It depends on the significance level required. And that, in turn, will depend on the cost of making the wrong decision. For ordinary use, a 95% significance level will require 1.96 sd

How many 4 letter codes are there if letters are not repeated?

If letters cannot be repeated, then there are 26 options for the first letter. There are then 25 options for the second letter. There are then 24 options for the third letter, and 23 for the fourth. This means that in total there are 26x25x24x23 options for the code. 26x25x24x23 = 358,800. Thus, there are 358,800 possible 4 letter codes.

What does the standard deviation of 2.686 means if the Mean is 5.96 and the number is 699?

The observation is more than 250 standard deviations (SD) away from the mean. For a normal distribution, the probability of being more than 3 SD from the mean is 0.0027 so the probability of an observation being 250 SD from the mean is infinitesimally small.

What is the probability of rolling a number greater than ten when 2 6 sided dice are thrown?

Well, you have a 6-4, or 4-6 and a 5-5, that's the only rolls that will role a ten. Out of 6*6=36 possible rolls. so, 3/36.

What is the probability of rolling a sum of 5 or 6 when you roll two number cubes?

There are 36 permutations of two dice. Of those, 9 add up to a sum of 5 or 6...

1 + 4 = 5

2 + 3 = 5

3 + 2 = 5

4 + 1 = 5

1 + 5 = 6

2 + 4 = 6

3 + 3 = 6

4 + 2 = 6

5 + 1 = 6

The probability, then, of rolling a 5 or 6 with two dice is 9 in 36, or 1 in 4, or 0.25.

What is the probability of rolling a 12 in a single roll of two dice?

The probability of rolling a sum of 12 in a single roll of two dice is 1 in 36, or about 0.0278.

How do you calculate sampling distribution parameters?

The answer depends on which parameters are to be calculated.

What is the probability of rolling a four on a die?

On a standard 6 sided die the probability of rolling four is 1/6

This is because there are 6 different numbers on it, and when you roll it you read only the top number. There are 6 to choose from and all 6 are equally probable, so that probability is 1 out of 6 = 0.1666....(a recurring 6) = 16.67% to 4 significant figures

How many ways can a hand of five cards consisting of four cards from one suit and one card from another suit be drawn from a standard deck of cards I think combination?

The probability of five cards being four cards from one suit and one card from another suit is the same as the probability of drawing four cards from one suit multiplied by the probability of drawing one card from another suit, multiplied by 5 (for each of the possible positions this other card can be drawn in).

The probability of drawing four cards from one suit is 12/51 x 11/50 x 10/49. The probability of drawing a fifth card from another suit is 39/48. All these numbers multiplied together (and multiplied by 5) come to 0.0429. So the probability of drawing a hand of five cards with four cards from one suit and one card from another is 5.29%

What are the odds of getting a sum divisible by three by rolling 2 dice?

If you have 12 possible numbers with multiple combinations then you should start out with making all the possible combinations; you will find theyre 20. Theyre four numbers out of the twleve that can be divisible by three; 3, 6, 9, and 12. There are 7 combinations where the combinations can equal those four numbers. So the odds of getting a sum divisible by three is 7/20.

What is the impact of the new point on the standard deviation?

The answer depends on the value of the new point. If the new value is near the mean then the new standard deviation (SD) will be smaller, if it is far away, the new SD will be larger.

Trending Questions
In the Yugioh trading card game if you play a card that lets you get a card from your deck do you shuffle your deck? What diagonals are always perpendicular? Having children the first year of marriage will cause the probability of divorce to double or triple? 7 red balls 4 white balls 7 orange balls what is the probability of picking the same color 2 out of 3 times? If A and B are independent events with probabilities 2 to 5 and 3 to 5 what is the probability that events A and B both occur? Probability word meaning 100 percent likely? When a number cube is rolled once what is the probability of getting an odd number greater than five? How many spades or kings in deck of 52 cards? If there is a 70 chance of rain tomorrow and a 50 chance of wind and rain what is the probability that it is windy given that it is rainy Round your answer to be the nearest percent? A fellow classmate tosses 3 coins and finds that 2 of them come up tails Which of the following is the best conclusion for her to come to? Which of these groups of relative frequencies would be best represented by a pie chart? What is the shortest way to figure out how many ways you can stack ten books on a shelf? How many 4 letter permutations can be formed from the word decagon? The likelihood that a event will happen? A bag of 6 marbles has numbers 1 through 6. If you pick 2 with out replacing what is the probability that the product is a multiple of 4? How do you calculate average sick time used per employee for a given year? How many necklace of 12 beads each can be made from 18 beads of different colours? What does it mean for an event to have a probability of 1? How many red face cards in a deck of cards? What is the probability of a vowel being picked at randoms from the word probability?